OpenStudy (anonymous):

how do you solve this (3x1/2x8) to the power of -2

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (mertsj):

\[(\frac{3x}{2x^8})^{-2}=(\frac{2x^8}{3x})^2=(\frac{2x^7}{3})^2=\frac{(2x^7)^2}{(3)^2}\]

OpenStudy (mertsj):

Can you finish it?